A 7200 square feet rectangular land need to be fenced
What is the length of the fence in yards if the length of the land is twice the width?

Answer by lwsshak3(11628)

A 7200 square feet rectangular land need to be fenced
What is the length of the fence in yards if the length of the land is twice the width?
***
let w=width
length=2w
length*width=area
2w*w=7200
2w^2=7200
w^2=7200/2=3600
w=60
length=2w=120ft or 120/3=40
length of the fence=40yds